# LED: LLM Enhanced Open-Vocabulary Object Detection without Human Curated Data Generation

### v4

@@ 1 Introduction
We argue that the latent representations of large multimodal language models already encode rich visual–linguistic concepts that need not be materialised into explicit labels. Building on this observation, we introduce L ED , an Latent Enhanced Open-Vocabulary Object Detection that realises a novel knowledge-fusion paradigm: instead of learning a detector from scratch or augmenting training with generated data, we directly fuse intermediate hidden states from a frozen MLLM into popular mainstream detector via a lightweight adapter. Specifically, as shown in Figure 1(a), existing approaches primarily distill LLM knowledge into the object detector via explicit data generation—such as synthesizing negative examples to mitigate hallucinations [38, 42, 29, 37, 47], or creating detailed annotations of spatial and logical relationships [16]. However, these explicit generation methods inherently constrain model robustness and generalizability. In contrast, as illustrated in Figure 1(b), our method directly fuses the high-dimensional semantic knowledge encapsulated within LLM hidden states into the detector, thereby preserving richer pretrained knowledge without explicit data augmentation.
@@ 2 Relate Work
Multi-Modal Large Language Models: Recent MLLMs have driven visual–language integration through modality alignment. CLIP pioneered contrastive learning on large-scale image–text pairs [27], and BLIP enhanced alignment with curated data [14, 15]. LLaVA applies GPT-4–generated multimodal instruction tuning to boost reasoning [20]. InternVL’s progressive alignment scales vision–language models to SOTA in foundation models [4, 5, 6, 9, 32]. LLaMA Adapter uses a zero-initialized attention mechanism with just 1.2 M parameters for prompt-based tuning, matching full fine-tuning [44]. L ED instead fuses high-level MLLM embeddings directly into downstream detectors.
@@ 3.1 Cross-modal Alignment
Stage 3 - Adapter Training: With the MLLM and detector frozen, we train the grounding decoder adapter (Block 3) and the MLP layer and LLM in the MLLM to integrate semantic cues from its intermediate features, thereby enabling the LLM to better focus on grounding tasks when processing visual tokens.
@@ 4.1 Dataset and Implementation Details
Stage 3 - Adapter training: We employ Grounding DINO [21] pretrain on Object365, GoldG and Cap4M as a unified detection pre-training framework across three datasets: Objects365 [31], COCO2017, and Flickr30k [39]. We implement our framework using InternVL2-1B (Qwen2-0.5B backbone) as the MLLM [4, 32, 9, 5, 6] and Open-GroundingDINO [21] as the detection codebase. ALL experiment configuration As shown in Section A.3 (see Supplement)
@@ 4.2.3 Different LLMs
OmniLabel | AP descr,categ | AP categ | AP descr | AP descr-pos |

### v5

@@ 1 Introduction
We argue that the latent representations of large multimodal language models already encode rich visual–linguistic concepts that need not be materialised into explicit labels. Building on this observation, we introduce L ED , an LLM Enhanced Open-Vocabulary Object Detection that realises a novel knowledge-fusion paradigm: instead of learning a detector from scratch or augmenting training with generated data, we directly fuse intermediate hidden states from a frozen MLLM into popular mainstream detector via a lightweight adapter. Specifically, as shown in Figure 1(a), existing approaches primarily distill LLM knowledge into the object detector via explicit data generation—such as synthesizing negative examples to mitigate hallucinations [38, 42, 29, 37, 47], or creating detailed annotations of spatial and logical relationships [16]. However, these explicit generation methods inherently constrain model robustness and generalizability. In contrast, as illustrated in Figure 1(b), our method directly fuses the high-dimensional semantic knowledge encapsulated within LLM hidden states into the detector, thereby preserving richer pretrained knowledge without explicit data augmentation.
@@ 2 Relate Work
Multi-Modal Large Language Models: Recent MLLMs have driven visual–language integration through modality alignment [11, 12]. CLIP pioneered contrastive learning on large-scale image–text pairs [27], and BLIP enhanced alignment with curated data [14, 15]. LLaVA applies GPT-4–generated multimodal instruction tuning to boost reasoning [20]. InternVL’s progressive alignment scales vision–language models to SOTA in foundation models [4, 5, 6, 9, 32]. LLaMA Adapter uses a zero-initialized attention mechanism with just 1.2 M parameters for prompt-based tuning, matching full fine-tuning [44]. L ED instead fuses high-level MLLM embeddings directly into downstream detectors.
@@ 3.1 Cross-modal Alignment
Stage 3 - Adapter Training: With the MLLM and detector frozen, we train the grounding decoder adapter (Block 3) and the MLP layer in the MLLM to integrate semantic cues from its intermediate features, thereby enabling the LLM to better focus on grounding tasks when processing visual tokens.
@@ 4.1 Dataset and Implementation Details
Stage 3 - Adapter training: We employ Grounding DINO [21] pretrain on Object365, GoldG and Cap4M as a unified detection pre-training framework across three datasets: Objects365 [31], COCO2017, and Flickr30k [39]. We implement our framework using InternVL2-1B (Qwen2-0.5B backbone) as the MLLM [4, 32, 9, 5, 6] and Open-GroundingDINO [21] as the detection codebase. All experiment configurations as shown in Section A.3 (see Supplement)
@@ 4.2.3 Different LLMs
Framework | AP descr,categ | AP categ | AP descr | AP descr-pos |